# On fixed points of pseudo-Anosov maps

### v1

@@ 1. Introduction
Let $S$ be an orientable finite-type surface with $\chi(S)<0$, and let $f\colon S\to S$ be a pseudo-Anosov homeomorphism of $S$. In this paper, we address the question of how combinatorial and geometric data associated to $f$ determines its number of fixed points.
@@
The notation $a\asymp b$ means that the two quantities are equal up to additive and multiplicative error depending only on $S$. See Section 2.1 for more details.
Theorem 1.1 fits into the following context. First, if we denote the stretch factor of $f$ by $\lambda$ and note that $\ell_{\mathcal{T}}(f)=\log\lambda$, then a theorem of Thurston [46] says that ${\rm Fix}(f^{n})\sim\lambda^{n}$. So the novelty of the theorem is that it counts fixed points of $f$ rather than fixed points of sufficiently large powers. Second, Theorem 1.1 asserts an upper and a lower bound on $\operatorname{Fix}(f)$. The upper bound is not new; indeed, the stronger bound $\operatorname{Fix}(f)\prec\lambda=\exp(\ell_{\mathcal{T}}(f))$ follows by considering a bounded-size Markov partition of $f$. See Remark 7.7 for details. So the real content of Theorem 1.1 is the lower bound on $\operatorname{Fix}(f)$, which is entirely new.
@@
To prove the claimed coarse inequality, it suffices to assume $\ell_{\mathcal{C}(S)}(\varphi)\geq 2$. This implies $\varphi=[f]$ is pseudo-Anosov, as reducible and periodic mapping classes have stable translation length $\ell_{\mathcal{C}(S)}(\varphi)=0$. It also implies $f$ is strongly irreducible. Since the systole map $\mathcal{T}(S)\to\mathcal{C}(S)$ is coarsely Lipschitz [26], the result follows from Theorem 1.1. ∎
@@
From a very different perspective, Heegaard Floer theory is a powerful package of homological invariants associated to closed oriented 33–manifolds and knots in 33–manifolds. To a closed oriented 33–manifold $Y$, the theory associates a finitely generated abelian group $\widehat{HF}(Y)$ that splits as a direct sum over classes in $H^{2}(Y)$. (There are also more refined invariants $HF^{+}(Y)$, $HF^{-}(Y)$, and $HF^{\infty}(Y)$, which are beyond the scope of our discussion.) To a null-homologous knot $K\subset Y$, it associates a finitely generated abelian group $\widehat{HFK}(Y,K)$ that splits as a direct sum over gradings valued in $\mathbb{Z}$. See Ozsváth and Szabó [36, 35, 34] and Rasmussen [43] for the original papers defining these invariants, and [37] for an excellent survey.
@@
(1.2) $\operatorname{Fix}(f)\leq 2\cdot\text{rank}_{\mathbb{Z}}\widehat{HF}(M_{f}).$
@@
(1.3) $\operatorname{Fix}(f)+1=\text{rank}\widehat{HFK}(Y,K;g(S)-1).$
@@
$\operatorname{Vol}(M_{f})\prec\log\operatorname{rank}_{\mathbb{Z}}\widehat{HF}(M_{f}),$
@@
$\leq\log\text{rank}_{\mathbb{Z}}\widehat{HF}(M_{f})+\log 2.$
@@
$\operatorname{Vol}(Y\smallsetminus K)\prec\log\operatorname{rank}\widehat{HFK}(Y,K),$
@@ 2.1. Coarse equalities, distance formulas, and cutoffs
Fix a surface $S$. For quantities $a$ and $b$ and a constant $C\geq 1$, we write $a\prec_{C}b$ to mean $a\leq C\cdot b+C$”. The equivalence relation $a\asymp_{C}b$ means $a\prec_{C}b$ and $b\prec_{C}a$”. As a common shorthand, we will write $a\asymp b$ to mean that the implicit constant $C$ depends only on the surface $S$, whereas $a,b$ depend on other parameters.
@@ 2.2. Arc and curve graphs
Let $S$ be an orientable surface of finite type, potentially with non-empty boundary or punctures. The arc and curve graph $\mathcal{AC}(S)$ is the graph whose vertices are isotopy classes of essential simple closed curves and simple proper arcs in $S$. Here, essential means that the curve or arc is not isotopic into a small neighborhood of a point, a boundary component, or a puncture. Two vertices are joined by an edge in $\mathcal{AC}(S)$ if they have disjoint representatives. If we follow the same construction with vertices restricted to be closed curves on $S$, we obtain the curve graph $\mathcal{C}(S)\subset\mathcal{AC}(S)$, and similarly restricting to arcs yields the arc graph $\mathcal{A}(S)\subset\mathcal{AC}(S)$.
@@ 2.4. Pseudo-Anosov homeomorphisms
A homeomorphism $f\colon S\to S$ is called pseudo-Anosov if there is a pair of transverse singular measured foliations $\mathcal{F}^{s},\mathcal{F}^{u}$ on $S$ and $\lambda>1$, called its stretch factor, so that $f(\mathcal{F}^{u})=\lambda\cdot\mathcal{F}^{u}$ and $f(\mathcal{F}^{s})=\frac{1}{\lambda}\cdot\mathcal{F}^{s}$. The measured foliations $\mathcal{F}^{s/u}$ are called the stable/unstable foliations of $f$.
Given these foliations (which are unique up to scaling their respective measures), there is a unique singular flat structure $X$ on $S$ whose singularities are precisely those of $\mathcal{F}^{s/u}$, and whose vertical and horizontal foliations are $\mathcal{F}^{s}$ and $\mathcal{F}^{u}$, respectively. Hence, our convention is that $f$ stretches along the horizontal leaves and contracts along the vertical leaves of $X$. More formally, $X$ can be described as a marked complex structure on $S$ together with a quadratic differential $q$ so that the image of $X$ under $f$ is equal to the image of $q$ under the time $\log(\lambda)$ Teichmüller flow.
Throughout the paper, it will often be convenient to puncture $S$ at the singularities of $f$. The resulting surface is denoted $\mathring{S}$ and there is an induced homeomorphism $\mathring{f}\colon\mathring{S}\to\mathring{S}$.
@@ 2.5. Veering triangulations
An (immersed) rectangle in $X$ is the image of an embedded rectangle in the universal cover $\widetilde{X}$, with its induced singular flat structure. By a rectangle in $\widetilde{X}$ we mean the embedded image of a (necessarily singularity-free) Euclidean rectangle whose vertical sides map to segments in $\widetilde{\mathcal{F}}^{s}$ and whose horizontal sides map to segments of $\widetilde{\mathcal{F}}^{u}$. A saddle connection in $X$ or $\widetilde{X}$ is a straight segment between two singularities, i.e. a geodesic in the flat metric that meets the singularities exactly at its endpoints. Particularly important to this paper are the saddle connections that span rectangles; these will be called either singularity-free saddle connections or veering edges, depending on the perspective. In slightly more detail, these are the saddle connections that arise by taking a rectangle $\widetilde{R}$ in $\widetilde{X}$ that contains two singularities in its boundary and connecting them with the straight line segment to produce a saddle connection $\sigma$. The induced (sub)-rectangle of $\widetilde{R}$ whose opposite (singular) corners lie at the endpoints $\sigma$ is called the spanning rectangle $R_{\sigma}$ of $\sigma$. The same terminology applies in $X$ after applying the covering projection (although spanning rectangles in $X$ are only immersed; see Lemmas 3.1 and 3.2).
@@ 3. The fundamental lemma
$\iota(\ell_{-},\sigma)\geq k\quad\text{hence}\quad d_{A}(\lambda^{-},\sigma)\geq k+1.$
@@ 5. Isolated pockets and annular-avoiding sections
By Lemma 4.1, $\mathring{W}$ is $\tau$–compatible. Let $T_{0}=T^{-}(\partial_{\tau}\mathring{W}),T_{1},\ldots,T^{n}=T^{-}(\partial_{\tau}\mathring{W})$ be an upward flip sequence from $T_{0}=T^{-}(\partial_{\tau}\mathring{W})$ to $T_{n}=T^{-}(\partial_{\tau}\mathring{W})$. Let $a$ be the largest integer in $\{0,\ldots,n\}$ such that $d_{\mathring{W}}(e,\lambda^{-})<6$ for some edge $e\subset T_{a-1}$ that projects to $\mathring{W}$; such an $a$ exists by Lemma 4.6. Similarly, let $b$ be the smallest integer in the sequence such that $d_{\mathring{W}}(e,\lambda^{+})<6$ for some edge $e\subset T_{b+1}$ projecting to $\mathring{W}$. The hypothesis $d_{\mathring{W}}(\lambda^{-},\lambda^{+})>12$ implies that $a<b$, and that $T_{a},T_{b}$ cannot share any edges over $\mathring{W}$. Now, let $V_{\mathring{W}}$ be the pocket over $\mathring{W}$ whose bottom side is contained in $T_{a}$ and whose top side is contained in $T_{b}$. By construction, every edge in $T_{a},\ldots,T_{b}$ that lies over $\mathring{W}$ satisfies $d_{\mathring{W}}(e,\lambda^{-})\geq 6$ and $d_{\mathring{W}}(e,\lambda^{+})\geq 6$. Thus $V_{\mathring{W}}$ is an isolated pocket, and property (1) also follows.
@@ 6. Distance formulas
$\log\#\operatorname{Fix}(f)=\log i(\sigma,f(\sigma))+O(1).$
@@
The second inequality holds because there are at most $|9\chi(S)|$ edges in $T$, and $\sigma$ was the chosen to maximize intersection numbers.
Now, consider the other direction. Since $T$ is annular-avoiding, Lemma 5.4 and Lemma 3.2 imply that the immersed spanning rectangle for $\sigma$ has degree uniformly bounded by $15$. Thus, by the fundamental lemma,
$\operatorname{Fix}(f)\geq\frac{1}{15}i(\sigma,f(\sigma)).\qed$
@@ 7.1. Remarks on the argument
$\log\#\operatorname{Fix}(f)\leq\log\lambda(f)+O(1)=\ell_{\mathcal{T}(S)}(f)+O(1).$
@@ 8.1. Applying distance shuffling
We will next apply Proposition 8.3 (distance shuffling) to prove Theorem 8.1. To begin the argument, recall that $\mathcal{D}_{f}$ is the set of subsurfaces that are overlapped by $f$. In parallel with this, define $\mathring{\mathcal{D}}_{f}$ to be the set of essential subsurfaces $\mathring{W}\subset\mathring{S}$ that are overlapped by $\mathring{f}$. The following lemma shows that (under a mild distance hypothesis) the subsurfaces in $\mathring{\mathcal{D}}_{f}$ are exactly the ones that fill to elements of $\mathcal{D}_{f}$.
@@
Let $T_{\sigma}$ be a section containing $\sigma$. Since $d_{\mathring{W}}(T_{\sigma},f(T_{\sigma}))\geq 10$, Theorem 4.7.(3) implies $\mathring{W}$ is $\tau$–compatible. Thus $\mathring{W}$ has an open representative $\mathrm{int}_{\tau}\mathring{W}$ whose boundary $\partial_{\tau}\mathring{W}$ consists of veering edges.
@@
By Lemma 8.6, any subsurface $\mathring{W}\subset\mathring{S}$ where $\big{[}d_{\mathring{W}}(\sigma,f(\sigma))\big{]}_{10}>0$ must actually belong to $\mathring{\mathcal{D}_{f}}$, hence $\mathcal{S}(\sigma,f(\sigma))\prec\mathcal{S}_{\mathring{\mathcal{D}_{f}}}(\sigma,f(\sigma))$. The reverse inequality is immediate, because every summand of $\mathcal{S}_{\mathring{\mathcal{D}_{f}}}(\sigma,f(\sigma))$ appears as a summand of $\mathcal{S}(\sigma,f(\sigma))$ by definition. ∎
We can now proceed to the proof of Theorem 8.1. The proof follows the same outline as the argument in Section 7. In particular, the coarse equality of Theorem 8.1 is proved by chaining together three coarse comparisons: Proposition 8.8 (which is a replacement for Proposition 7.1), Proposition 8.9 (which functions as a replacement for Proposition 7.4), and Proposition 7.5 (used in its original form).

### v2

@@ 1. Introduction
Let $S$ be an orientable finite-type surface without boundary, and with $\chi(S)<0$. Consider a pseudo-Anosov homeomorphism $f\colon S\to S$ (see Section 2.4 for the definition). In this paper, we address the question of how combinatorial and geometric data associated to $f$ determines its number of fixed points.
@@
Since the usual definition of a pseudo-Anosov map (Section 2.4) assumes that the domain of $f$ can have punctures but no boundary, this is our assumption throughout the paper. The notation $a\asymp b$ means that the two quantities are equal up to additive and multiplicative error depending only on $S$. See Section 2.1 for more details.
Theorem 1.1 fits into the following context. First, if we denote the stretch factor of $f$ by $\lambda$ and note that $\ell_{\mathcal{T}}(f)=\log\lambda$, then a theorem of Thurston [46] says that ${\rm Fix}(f^{n})\sim\lambda^{n}$. So the novelty of the theorem is that it counts fixed points of $f$ rather than fixed points of sufficiently large powers. Second, Theorem 1.1 asserts an upper and a lower bound on $\operatorname{Fix}(f)$. The upper bound is not new; indeed, the stronger bound $\log\#\operatorname{Fix}(f)\leq\ell_{\mathcal{T}}(f)+O(\log|\chi(S)|)$ follows by considering a bounded-size Markov partition of $f$. See Remark 7.7 for details. So the real content of Theorem 1.1 is the lower bound on $\operatorname{Fix}(f)$, which is entirely new.
@@
To prove the claimed coarse inequality, it suffices to assume $\ell_{\mathcal{C}(S)}(\varphi)>1$. This implies $\varphi=[f]$ is pseudo-Anosov, as reducible and periodic mapping classes have stable translation length $\ell_{\mathcal{C}(S)}(\varphi)=0$. It also implies $f$ is strongly irreducible. Since the systole map $\mathcal{T}(S)\to\mathcal{C}(S)$ is coarsely Lipschitz [26], the result follows from Theorem 1.1. ∎
@@
From a very different perspective, Heegaard Floer theory is a powerful package of homological invariants associated to closed oriented 33–manifolds and knots in 33–manifolds. To a closed oriented 33–manifold $Y$, the theory associates a finitely generated abelian group $\widehat{\operatorname{HF}}(Y)$ that splits as a direct sum over classes in $H^{2}(Y)$. (There are also more refined invariants $\operatorname{HF}^{+}(Y)$, $\operatorname{HF}^{-}(Y)$, and $\operatorname{HF}^{\infty}(Y)$, which are beyond the scope of our discussion.) To a null-homologous knot $K\subset Y$, it associates a finitely generated abelian group $\widehat{\operatorname{HFK}}(Y,K)$ that splits as a direct sum over gradings valued in $\mathbb{Z}$. See Ozsváth and Szabó [36, 35, 34] and Rasmussen [43] for the original papers defining these invariants, and [37] for an excellent survey.
@@
(1.2) $\operatorname{Fix}(f)\leq 2\cdot\text{rank}_{\mathbb{Z}}\,\widehat{\operatorname{HF}}(M_{f}).$
@@
(1.3) $\operatorname{Fix}(f)+1=\text{rank}\,\widehat{\operatorname{HFK}}(Y,K;g(S)-1).$
@@
$\operatorname{Vol}(M_{f})\prec\log\operatorname{rank}_{\mathbb{Z}}\widehat{\operatorname{HF}}(M_{f}),$
@@
$\leq\log\text{rank}_{\mathbb{Z}}\,\widehat{\operatorname{HF}}(M_{f})+\log 2.$
@@
$\operatorname{Vol}(Y\smallsetminus K)\prec\log\operatorname{rank}\,\widehat{\operatorname{HFK}}(Y,K),$
@@ 2.1. Coarse equalities, distance formulas, and cutoffs
Fix a surface $S$. For quantities $a$ and $b$ and a constant $C\geq 1$, we write $a\prec_{C}b$ to mean $a\leq C\cdot b+C$”. The equivalence relation $a\asymp_{C}b$ means $a\prec_{C}b$ and $b\prec_{C}a$”. As a common shorthand, we will write $a\prec b$ or $a\asymp b$ to mean that the implicit constant $C$ depends only on the surface $S$, whereas $a,b$ depend on other parameters. The relation $\prec$ is called coarse inequality, and $\asymp$ is called coarse equality.
@@ 2.2. Arc and curve graphs
Let $S$ be an orientable surface of finite type, potentially with non-empty boundary or punctures. An arc in $S$ is a proper embedding of an interval (possibly open, closed, or neither), where “proper” means that an endpoint of the interval is mapped into $\partial S$ and the preimage of a compact set is compact. Less formally, this latter condition means that each open end of the interval is mapped to an end of $S$. By an isotopy of an arc, we mean an isotopy through such proper embeddings.
The arc and curve graph $\mathcal{AC}(S)$ is the graph whose vertices are isotopy classes of essential simple closed curves and essential arcs in $S$. Here, essential means that the curve or arc is not isotopic into a small neighborhood of a point, a boundary component, or a puncture. Two vertices are joined by an edge in $\mathcal{AC}(S)$ if they have disjoint representatives. If we follow the same construction with vertices restricted to be closed curves on $S$, we obtain the curve graph $\mathcal{C}(S)\subset\mathcal{AC}(S)$, and similarly restricting to arcs yields the arc graph $\mathcal{A}(S)\subset\mathcal{AC}(S)$.
@@ 2.4. Pseudo-Anosov homeomorphisms
Let $S$ be a surface without boundary, possibly with punctures. A homeomorphism $f\colon S\to S$ is called pseudo-Anosov if there is a pair of transverse singular measured foliations $\mathcal{F}^{s},\mathcal{F}^{u}$ on $S$ and a constant $\lambda>1$, called its stretch factor, so that $f(\mathcal{F}^{u})=\lambda\cdot\mathcal{F}^{u}$ and $f(\mathcal{F}^{s})=\frac{1}{\lambda}\cdot\mathcal{F}^{s}$. The measured foliations $\mathcal{F}^{s/u}$ are called the stable/unstable foliations of $f$.
Given the singular measured foliations $\mathcal{F}^{s}$ and $\mathcal{F}^{u}$, there is a unique singular flat structure $X$ on $S$ whose singularities are precisely those of $\mathcal{F}^{s/u}$, and whose whose vertical and horizontal foliations are $\mathcal{F}^{s}$ and $\mathcal{F}^{u}$, respectively. Since the measures on $\mathcal{F}^{s/u}$ are unique up to scaling, the singular flat structure $X$ is also unique up to horizontal and vertical rescaling. In particular, our convention is that $f$ stretches along the horizontal leaves and contracts along the vertical leaves of $X$. More formally, $X$ can be described as a marked complex structure on $S$ together with a quadratic differential $q$ so that the image of $X$ under $f$ is equal to the image of $q$ under the time $\log(\lambda)$ Teichmüller flow.
Throughout the paper, it will often be convenient to puncture $S$ at the singularities of $f$. The resulting surface is denoted $\mathring{S}$ and there is an induced pseudo-Anosov homeomorphism $\mathring{f}\colon\mathring{S}\to\mathring{S}$.
@@ 2.5. Veering triangulations
We continue to use the notation of Section 2.4, where $X$ is a singular flat structure associated to the pseudo-Anosov map $f$. An (immersed) rectangle in $X$ is the image of an embedded rectangle in the universal cover $\widetilde{X}$, with its induced singular flat structure. By a rectangle in $\widetilde{X}$ we mean the embedded image of a (necessarily singularity-free) Euclidean rectangle whose vertical sides map to segments in $\widetilde{\mathcal{F}}^{s}$ and whose horizontal sides map to segments of $\widetilde{\mathcal{F}}^{u}$. A saddle connection in $X$ or $\widetilde{X}$ is a straight segment between two singularities, i.e. a geodesic in the flat metric that meets the singularities exactly at its endpoints. Particularly important to this paper are the saddle connections that span rectangles; these will be called either singularity-free saddle connections or veering edges, depending on the perspective. In slightly more detail, these are the saddle connections that arise by taking a rectangle $\widetilde{R}$ in $\widetilde{X}$ that contains two singularities in its boundary and connecting them with the straight line segment to produce a saddle connection $\sigma$. The induced (sub)-rectangle of $\widetilde{R}$ whose opposite (singular) corners lie at the endpoints $\sigma$ is called the spanning rectangle $R_{\sigma}$ of $\sigma$. The same terminology applies in $X$ after applying the covering projection (although spanning rectangles in $X$ are only immersed; see Lemmas 3.1 and 3.2).
@@ 3. The fundamental lemma
$i(\ell_{-},\sigma)\geq k\quad\text{hence}\quad d_{A}(\lambda^{-},\sigma)\geq k+1.$
@@ 5. Isolated pockets and annular-avoiding sections
By Lemma 4.1, $\mathring{W}$ is $\tau$–compatible. Let $T_{0}=T^{-}(\partial_{\tau}\mathring{W}),T_{1},\ldots,T_{n}=T^{+}(\partial_{\tau}\mathring{W})$ be an upward flip sequence from $T_{0}=T^{-}(\partial_{\tau}\mathring{W})$ to $T_{n}=T^{+}(\partial_{\tau}\mathring{W})$. Let $a$ be the largest integer in $\{0,\ldots,n\}$ such that $d_{\mathring{W}}(e,\lambda^{-})<6$ for some edge $e\subset T_{a-1}$ that projects to $\mathring{W}$; such an $a$ exists by Lemma 4.6. Similarly, let $b$ be the smallest integer in the sequence such that $d_{\mathring{W}}(e,\lambda^{+})<6$ for some edge $e\subset T_{b+1}$ projecting to $\mathring{W}$. The hypothesis $d_{\mathring{W}}(\lambda^{-},\lambda^{+})>12$ implies that $a<b$, and that $T_{a},T_{b}$ cannot share any edges over $\mathring{W}$. Now, let $V_{\mathring{W}}$ be the pocket over $\mathring{W}$ whose bottom side is contained in $T_{a}$ and whose top side is contained in $T_{b}$. By construction, every edge in $T_{a},\ldots,T_{b}$ that lies over $\mathring{W}$ satisfies $d_{\mathring{W}}(e,\lambda^{-})\geq 6$ and $d_{\mathring{W}}(e,\lambda^{+})\geq 6$. Thus $V_{\mathring{W}}$ is an isolated pocket, and property (1) also follows.
@@ 6. Distance formulas
$\log\#\operatorname{Fix}(f)=\log i(\sigma,f(\sigma))+O(\log|\chi(S)|).$
@@
The second inequality holds because there are at most $|9\chi(S)|$ edges in $T$, and $\sigma$ was chosen to maximize intersection numbers.
Now, consider the other direction. Since $T$ is annular-avoiding, Lemma 5.4 and Lemma 3.2 imply that the immersed spanning rectangle for $\sigma$ has degree uniformly bounded by $13$. Thus, by the fundamental lemma,
$\operatorname{Fix}(f)\geq\frac{1}{13}i(\sigma,f(\sigma)).\qed$
@@ 7.1. Remarks on the argument
$\log\#\operatorname{Fix}(f)\leq\log\lambda(f)+O(\log|\chi(S)|)=\ell_{\mathcal{T}(S)}(f)+O(\log|\chi(S)|).$
@@ 8. Without strong irreducibility
Proposition 8.3 (distance shuffling) will be proved in Section 8.2.
@@ 8.1. Applying distance shuffling
For now, we will apply Proposition 8.3 to prove Theorem 8.1. To begin the argument, recall that $\mathcal{D}_{f}$ is the set of subsurfaces that are overlapped by $f$. In parallel with this, define $\mathring{\mathcal{D}}_{f}$ to be the set of essential subsurfaces $\mathring{W}\subset\mathring{S}$ that are overlapped by $\mathring{f}$. The following lemma shows that (under a mild distance hypothesis) the subsurfaces in $\mathring{\mathcal{D}}_{f}$ are exactly the ones that fill to elements of $\mathcal{D}_{f}$.
@@
Let $T_{\sigma}$ be a section containing $\sigma$. Since $d_{\mathring{W}}(T_{\sigma},f(T_{\sigma}))\geq 10$, Theorem 4.7(3) implies $\mathring{W}$ is $\tau$–compatible. Thus $\mathring{W}$ has an open representative $\mathrm{int}_{\tau}\mathring{W}$ whose boundary $\partial_{\tau}\mathring{W}$ consists of veering edges.
@@
By Lemma 8.6, any subsurface $\mathring{W}\subset\mathring{S}$ where $\big[d_{\mathring{W}}(\sigma,f(\sigma))\big]_{10}>0$ must actually belong to $\mathring{\mathcal{D}_{f}}$, hence $\mathcal{S}(\sigma,f(\sigma))\prec\mathcal{S}_{\mathring{\mathcal{D}_{f}}}(\sigma,f(\sigma))$. The reverse inequality is immediate, because every summand of $\mathcal{S}_{\mathring{\mathcal{D}_{f}}}(\sigma,f(\sigma))$ appears by definition as a summand of $\mathcal{S}(\sigma,f(\sigma))$. ∎
We can now proceed to the proof of Theorem 8.1. The proof follows the same outline as the argument in Section 7. In particular, the coarse equality of Theorem 8.1 is proved by chaining together three coarse comparisons: Proposition 8.8 (which is a replacement for Proposition 7.1), Proposition 8.9 (which is a replacement for Proposition 7.4), and Proposition 7.5 (used in its original form).